OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (amistre64):

what is the first term?

OpenStudy (anonymous):

f(0) = -5 ?

OpenStudy (amistre64):

yes, so they give you the first one what is the second term?

OpenStudy (amistre64):

we can narrow the options by eliminating the ones that have the wrong first and second terms, then the third term gets us to only one option that fits

OpenStudy (amistre64):

we can use the rule given to determine f(2) f(n) = -4•f(n -1) - 3•f(n - 2) f(2) = -4•f(2 -1) - 3•f(2 - 2) f(2) = -4•f(1) - 3•f(0) ^^ ^^ we already know these values to plug into the rule

OpenStudy (anonymous):

ok so what now ?

OpenStudy (amistre64):

that IS the "what now" you work it out

OpenStudy (amistre64):

or you ask questions about what it is you dont understand

OpenStudy (anonymous):

oh ok hold on

OpenStudy (anonymous):

i dont get what im suppose to work out ....

OpenStudy (amistre64):

f(n) = -4•f(n -1) - 3•f(n - 2) f(2) = -4•f(2 -1) - 3•f(2 - 2) f(2) = -4•f(1) - 3•f(0) ^^ ^^ we already know these values to plug into the rule

OpenStudy (amistre64):

f(0) = -5 f(1) = 20

OpenStudy (amistre64):

wherever you see f(0), replace it by -5 wherever you see f(1), replace it by 20

OpenStudy (anonymous):

what if i see f(2)

OpenStudy (amistre64):

f(2) is what we are calculating ... f(2) will equal ______________

OpenStudy (anonymous):

f(2) = -4•f(1) - 3•f(0) =-65 correct?

OpenStudy (amistre64):

very good :)

OpenStudy (amistre64):

so what we are looking for is a sequence that starts out: f(0), f(1), f(2) -5, 20, -65 only one of the options starts out like this

OpenStudy (anonymous):

the last one ..

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you very much for explaining

OpenStudy (amistre64):

good luck :)